OpenStudy (anonymous):

Mia is three years older than twice her sister Brooke’s age. The sum of their ages is less than 30. What is the greatest age Brooke could be? A. 7 B. 9 C. 8 D. 10 Question Resources

OpenStudy (anonymous):

M = 2B + 3 M + B < 30 M = 30 - B 30 - B = 2B + 3 3B = 27 B=?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Your welcome

OpenStudy (lieutenantgeneral):

D. 10 and Welcome.

OpenStudy (anonymous):

I failed

OpenStudy (anonymous):

Let Brooke be x years old Mia is 2x + 3 years old Their total age is x + (2x + 3) = 3x + 3 Their age is less than 30, so we have: 3x + 3 < 30 3x < 27 x <9 Do you get it now?

OpenStudy (anonymous):

yea but I still failed